# Constrained Synthesis with Projected Diffusion Models

### v2

@@ 1 Introduction
Generative diffusion models excel at robustly synthesizing content from raw noise through a sequential denoising process [23, 14]. They have revolutionized high-fidelity creation of complex data, and their applications have rapidly expanded beyond mere image synthesis, finding relevance in areas such as engineering [28, 31], automation [3, 16], chemistry [1, 15], and medical analysis [2, 6]. Although diffusion models excel at generating content that is coherent and aligns closely with the original data distribution, their direct application in scenarios requiring stringent adherence to predefined criteria poses significant challenges. Particularly the use of diffusion models in domains where the generated data needs to not only resemble real-world examples but also rigorously comply with established specifications and physical laws remains an open challenge.
@@
Contributions. In summary, this paper makes the following key contributions: (1) It introduces PDM, a new framework that augments diffusion-based synthesis with arbitrary constraints in order to generate content with high fidelity that also adheres to the imposed specifications. (2) Extensive experiments across various domains demonstrate PDM’s effectiveness. These include adherence to morphometric properties in real-world material science experiments, physics-informed motion governed by ordinary differential equations, trajectory optimization in motion planning, and constrained human motion synthesis, showcasing PDM’s ability to produce content that adheres to both complex constraints and physical principles. (3) We further show that PDM is able to generate out-of-distribution samples that meet stringent constraints, even in scenarios with extremely sparse training data and when the training data does not satisfy the required constraints. (4) Finally, we provide a theoretical basis elucidating the ability of PDM to generate highly accurate content while ensuring constraint compliance, underpinning the practical implications of this approach.
@@ 4 Constrained generative diffusion
The application of the reverse diffusion process of score-based models is characterized by iteratively fitting the initial noisy samples $\bm{x}_{T}$ to the learned approximation of $p(\bm{x}_{0})$. This process is akin to maximizing the density function $\log q(\bm{x}_{t}|\bm{x}_{0})$ at each given noise level, the gradients of which are estimated by $\mathbf{s}_{\theta}(\bm{x}_{t}^{i},t)$. In traditional score-based models, at any point throughout the reverse process, $\bm{x}_{t}$ is unconstrained. When these samples are required to satisfy some constraints, the objective remains unchanged, but the solution to this optimization must fall within a feasible region $\mathbf{C}$,
@@
Operationally, the negative log likelihood is minimized at each step of the reverse Markov chain, as the process transitions from $\bm{x}_{T}$ to $\bm{x}_{0}$. In this regard, and importantly, the objective of the PDM’s sampling process is aligned with that of traditional score-based diffusion models.
In an unconstrained setting, this optimization is formulated such that a variation of the traditional gradient ascent algorithm is used to iteratively transform a sample from the Gaussian distribution $q(\bm{x}_{T}|\bm{x}_{0})$ to a sample from the learned distribution $q(\bm{x}_{1}|\bm{x}_{0})$. The update step is provided by:
where $\bm{\epsilon}$ is standard normal and $\gamma_{t}>0$ is the step size. This step is repeated ${M}$ times for each $\bm{x}_{T}$ to $\bm{x}_{0}$. To prevent a deterministic behavior, an additional term is added to the gradient ascent algorithm, $\sqrt{2\gamma_{t}}\bm{\epsilon}$, drawing from Stochastic Gradient Langevin Dynamics [25]. To avoid low density data regions, the sample is optimized to conform to the previous distribution in the Markov chain before proceeding to the consecutive distribution, the transitions being ensured by setting $\bm{x}_{t-1}^{0}=\bm{x}_{t}^{M}$, where as $\bm{x}_{t}^{M}$ is the final iterate of the previous time step.
@@ 4.1 Constrained guidance through iterative projections
The score network $\mathbf{s}_{\theta}(\bm{x}_{t},t)$ directly estimates the first-order derivatives of eq. 3a, providing the necessary gradients for iterative gradient-based optimization defined in eq. 4. In the presence of constraints (3b), however, an alternative iterative method is necessary to guarantee feasibility. PDM models a projected guidance approach to provide this constraint-aware optimization process.
@@
where $\mathbf{C}$ is the set of constraints and $\mathcal{P}_{\mathbf{C}}$ is a projection onto $\mathbf{C}$. Hence, iteratively throughout the Markov chain, a gradient step is taken to minimize the objective defined by eq. 3a while ensuring feasibility. Convergence is guaranteed for convex constraints sets [20] and empirical evidence in §5 showcases the applicability of this methods to arbitrary constraint sets. Importantly, the projection operators can be warm-started during the repeated sampling step providing a piratical solution even for hard non-convex constrained regions. The full sampling process is detailed in Algorithm 1.
@@ 8      $\bm{x}_{t-1}^{0}\leftarrow\bm{x}_{t}^{M}$
Ground P(%) Generative Methods
PDM Cond Post+ Cond+
[…]
@@ 5.3 Constrained trajectories (nonconvex constraints)
The next experiment showcase the ability of PDM to handle nonconvex constraints. Path planning is a classic optimization problem which is integral to finding smooth, collision-free paths in autonomous systems. This setting consists of minimizing the path length while avoiding path intersection with various obstacles in a given topography. Recent research has demonstrated the use of diffusion models for these motion planning objectives [3]. In this task, the diffusion model predicts a series of points, $p_{0},p_{1},\ldots,p_{N}$, where each pair of consecutive points represents a line segment. The start and end points for this path are determined pseudo-randomly for each problem instance, with the topography remaining constant across different instances. Additionally, the problem presents obstacles at inference time (shown in red on Figure 6), that were not present during training, rendering a portion of the training data infeasible, and thus testing the generalization of these methods. The performance is evaluated on two sets of maps adapted from Carvalho et al., shown in Figure 6. The main challenge in this setting is the non-convex nature of the contraints.
Figure 5: Constrained trajecotires synthetized by PDM on two topographies (Tp1, left and Tp2, right).
@@ 5.4 Physics-informed motion (ODEs and out-of-distribution constraints)
6 Effectiveness of PDM: A theoretical justification
To shed light on the PDM ability to guide the generated samples to the constrained distribution, this section provides a theoretically grounded motivation for the use of iterative projections. The analysis assumes that the feasible region $\bm{C}$ is a convex set. We start by defining the update step.
Definition 6.1.
Theorem 6.2.
Corollary 6.3.
For arbitrary small $\xi>0$, there exist $t$ and $\bm{i}\geq\bar{I}$ such that:
Feasibility guarantees.
Corollary 6.4.

### v3

@@ 1 Introduction
Generative diffusion models excel at robustly synthesizing content from raw noise through a sequential denoising process [24, 14]. They have revolutionized high-fidelity creation of complex data, and their applications have rapidly expanded beyond mere image synthesis, finding relevance in areas such as engineering [29, 34], automation [3, 16], chemistry [1, 15], and medical analysis [2, 6]. However, although diffusion models excel at generating content that is coherent and aligns closely with the original data distribution, their direct application in scenarios requiring stringent adherence to predefined criteria poses significant challenges. Particularly the use of diffusion models in scientific and engineering domains where the generated data needs to not only resemble real-world examples but also rigorously comply with established specifications and physical laws remains an open challenge.
@@
Contributions. In summary, this paper makes the following key contributions: (1) It introduces PDM, a new framework that augments diffusion-based synthesis with arbitrary constraints in order to generate content with high fidelity that also adheres to the imposed specifications. The paper elucidates the theoretical foundation that connects the reverse diffusion process to an optimization problem, facilitating the direct incorporation of constraints into the reverse process of score-based diffusion models. (2) Extensive experiments across various domains demonstrate PDM’s effectiveness. These include adherence to morphometric properties in real-world material science experiments, physics-informed motion governed by ordinary differential equations, trajectory optimization in motion planning, and constrained human motion synthesis, showcasing PDM’s ability to produce content that adheres to both complex constraints and physical principles. (3) We further show that PDM is able to generate out-of-distribution samples that meet stringent constraints, even in scenarios with extremely sparse training data and when the training data does not satisfy the required constraints. (4) Finally, we provide a theoretical basis elucidating the ability of PDM to generate highly accurate content while ensuring constraint compliance, underpinning the practical implications of this approach.
@@ 4 Constrained generative diffusion
This section establishes a theoretical framework that connects the reverse diffusion process as an optimization problem. This perspective facilitates the incorporation of constraints directly into the process, resulting in the constrained optimization formulation presented in Equation (6).
The application of the reverse diffusion process of score-based models is characterized by iteratively transforming the initial noisy samples $\bm{x}_{T}$ back to a data sample $\bm{x}_{0}$ following the learned data distribution $q(\bm{x}_{0})$. This transformation is achieved by iteratively updating the sample using the estimated score function $\nabla_{\bm{x}_{t}}\log q(\bm{x}_{t}|\bm{x}_{0})$, where $q(\bm{x}_{t}|\bm{x}_{0})$ is the data distribution at time $t$. At each time step $t$, starting from $\bm{x}_{t}^{0}$, the process performs $M$ iterations of Stochastic Gradient Langevin Dynamics (SGLD) [30]:
where $\bm{\epsilon}$ is standard normal, $\gamma_{t}>0$ is the step size, and $\nabla_{\bm{x}_{t}^{i}}\log q(\bm{x}_{t}^{i}|\bm{x}_{0})$ is approximated by the learned score function $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$.
4.1 Casting the reverse process as an optimization problem
First note that SGLD is derived from discretizing the continuous-time Langevin dynamics, which are governed by the stochastic differential equation:
$d\bm{X}(t)=\nabla\log q(\bm{X}(t))\,dt+\sqrt{2}\,d\bm{B}(t),$ (4)
where $\bm{B}(t)$ is standard Brownian motion. Under appropriate conditions, the stationary distribution of this process is $q(\bm{x}_{t})$ [22], implying that samples generated by Langevin dynamics will, over time, be distributed according to $q(\bm{x}_{t})$. In practice, these dynamics are simulated using a discrete-time approximation, leading to the SGLD update in Equation (3). Therein the noise term $\sqrt{2\gamma_{t}}\,\bm{\epsilon}_{t}^{i}$ allows the algorithm to explore the probability landscape and avoid becoming trapped in local maxima.
Next notice that, as detailed in [31, 30], under some regularity conditions this iterative SGLD algorithm converges toward a stationary point, bounded by $\frac{d^{2}}{\sigma^{1/4}\lambda^{*}}\log(1/\epsilon),$ where, $\sigma^{2}$ represents the variance schedule, $\lambda^{*}$ denotes the uniform spectral gap of the Langevin diffusion, and $d$ is the dimensionality of the problem. Thus, as the reverse diffusion process progresses towards $T\to 0$, and the variance schedule decreases, the stochastic component becomes negligible, and SGLD transitions toward deterministic gradient ascent on $\log q(\bm{x}_{t})$. In the limit of vanishing noise, the update rule simplifies to:
$\bm{x}_{t}^{i+1}=\bm{x}_{t}^{i}+\gamma_{t}\nabla_{\bm{x}}\log q(\bm{x}_{t}^{i}|\bm{x}_{0}),$ (5)
which is standard gradient ascent aiming to maximize $\log q(\bm{x}_{t})$. This allow us to view the reverse diffusion process as an optimization problem minimizing the negative log-likelihood of the data distribution $q(\bm{x}_{t}|\bm{x}_{0})$ at each time step $t$.
In traditional score-based models, at any point throughout the reverse process, $\bm{x}_{t}$ is unconstrained. When these samples are required to satisfy some constraints, the objective remains unchanged, but the solution to this optimization must fall within a feasible region $\mathbf{C}$, and thus the optimization problem formulation becomes:
@@ 4.1 Casting the reverse process as an optimization problem
Operationally, the negative log likelihood is minimized at each step of the reverse Markov chain, as the process transitions from $\bm{x}_{T}$ to $\bm{x}_{0}$. In this regard, and importantly, the objective of the PDM’s reverse sampling process is aligned with that of traditional score-based diffusion models.
@@ 4.2 Constrained guidance through iterative projections
The score network $\mathbf{s}_{\theta}(\bm{x}_{t},t)$ directly estimates the first-order derivatives of Equation (6a), providing the necessary gradients for iterative gradient-based updates defined in Equation (3). In the presence of constraints (6b), however, an alternative iterative method is necessary to guarantee feasibility. PDM models a projected guidance approach to provide this constraint-aware optimization process.
@@
where $\mathbf{C}$ is the set of constraints and $\mathcal{P}_{\mathbf{C}}$ is a projection onto $\mathbf{C}$. Hence, iteratively throughout the Markov chain, a gradient step is taken to minimize the objective defined by Equation (6a) while ensuring feasibility. Convergence is guaranteed for convex constraints sets [20] and empirical evidence in §6 showcases the applicability of this methods to arbitrary constraint sets. Importantly, the projection
@@ 8      $\bm{x}_{t-1}^{0}\leftarrow\bm{x}_{t}^{M}$
operators can be warm-started during the repeated sampling step providing a piratical solution even for hard non-convex constrained regions. The full sampling process is detailed in Algorithm 1.
@@
5 Effectiveness of PDM: A theoretical justification
Next, we theoretically justify the use of iterative projections to guide the sample to the constrained distribution. The analysis assumes that the feasible region $\bm{C}$ is a convex set. All proofs are reported in the Appendix. We start by defining the update step.
Definition 5.1.
Theorem 5.2.
Corollary 5.3.
For arbitrary small $\xi>0$, there exist $t$ and $\bm{i}\geq\bar{I}$ such that:
Feasibility guarantees.
Corollary 5.4.
@@ 6.1 Constrained materials (low data regimes and constraint-violating distributions)
Ground P(%) Generative Methods
PDM Cond Post+ Cond+
[…]
@@ 6.3 Constrained trajectories (nonconvex constraints)
The next experiment showcase the ability of PDM to handle nonconvex constraints. Path planning is a classic optimization problem which is integral to finding smooth, collision-free paths in autonomous systems. This setting consists of minimizing the path length while avoiding path intersection with various obstacles in a given topography. Recent research has demonstrated the use of diffusion models for these motion planning objectives [3]. In this task, the diffusion model predicts a series of points, $p_{0},p_{1},\ldots,p_{N}$, where each pair of consecutive points represents a line segment. The start and end points for this path are determined pseudo-randomly for each problem instance, with the topography remaining constant across different instances. Additionally, the problem presents obstacles at inference time (shown in red on Figure 6), that were not present during training, rendering a portion of the training data infeasible, and thus testing the generalization of these methods. The performance is evaluated on two sets of maps adapted from Carvalho et al., shown in Figure 6. The main challenge in this setting is the non-convex nature of the constraints.
Figure 5: Constrained trajectories synthesized by PDM on two topographies (Tp1, left and Tp2, right).
@@ 8 Conclusions
9 Acknowledgments
This research is partially supported by NSF grants 2334936, 2334448, and NSF CAREER Award 2401285. Fioretto is also supported by an Amazon Research Award and a Google Research Scholar Award. The authors acknowledge Research Computing at the University of Virginia for providing computational resources that have contributed to the results reported within this paper. The views and conclusions of this work are those of the authors only.
Authors Contributions
JC and FF formulated the research question, designed the methodology, developed the theoretical analysis, and wrote the manuscript. Moreover, JC contributed to developing the code and performed the experimental analysis. SB acquired the data for the micro-structure experiment, formulated the desired properties for such experiment, and participated in the interpretation of the results.